OpenStudy (anonymous):

A 504 g mass oscillates with an amplitude of 11.0 cm on a spring whose spring constant is 20.0 N/m . Determine the period. Determine the maximum speed.

OpenStudy (michele_laino):

hint: the perios \(T\) is given by the subsequent formula: \[\huge T = 2\pi \sqrt {\frac{m}{K}} \] wherein \(K\) is the spring constant, and \(m\) is the oscillating mass. Whereas the maximum speed \(v_{max}\) is related to the amplitude \(A\), by means of this formula: \[\huge \frac{1}{2}mv_{\max }^2 = \frac{1}{2}K{A^2}\]

OpenStudy (michele_laino):

namely, it is the energy conservation law, which holds, if we can neglect any friction force

OpenStudy (whpalmer4):

for an additional test of your understanding, at what point is the maximum speed seen?